# Graph Neural Network-Based Reinforcement Learning for Controlling Biological Networks: The GATTACA Framework

### v1

@@ 2.3 Reinforcement learning
Nearly all RL algorithms involve estimating value functions. For MDPs, two such functions are commonly considered. First, the state-value function $v_{\pi}(s)$ defined as
$v_{\pi}(s)\triangleq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s]=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\bigm{|}S_{t}=s\right]\!,$
where $\mathbb{E}_{\pi}[\cdot]$ is the expected value of a random variable given that the agent follows policy $\pi$. Second, the action-value function $q_{\pi}(s,a)$ is defined as the expected return starting from $s$, taking the action $a$, and thereafter following policy $\pi$:
$q_{\pi}(s,a)$ $\triangleq\mathbb{E}_{\pi}[G_{t}\mid S_{t}=s,A_{t}=a]$
$=\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\biggm{|}S_{t}=s,A_{t}=a\right]\!.$
The value functions satisfy particular recursive relationships expressed in the form of Bellman equations, i.e. one for $v_{\pi}$:
$v_{\pi}(s)=\sum_{a\in\mathcal{A}}\pi(s,a)\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma v_{\pi}(s^{\prime})\Bigr{]}$ (1)
and one for $q_{\pi}$:
$q_{\pi}(s,a)=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma\sum_{a^{\prime}\in\mathcal{A}}\pi(s^{\prime},a^{\prime})q_{\pi}(s^{\prime},a^{\prime})\Bigr{]}.$ (2)
Value functions define a partial order over policies, that is, $\pi\geq\pi^{\prime}$ if and only if $v_{\pi}(s)\geq v_{\pi}^{\prime}(s)$ for all $s\in\mathcal{S}$. There always exists at least one policy which is better than or equal to all other policies [31]. Such policy is referred to as an optimal policy and all optimal policies are denoted by $\pi_{*}$. Optimal policies share the same optimal state-value function, denoted $v_{*}$, and defined as
$v_{*}(s)\triangleq\max_{\pi}v_{\pi}(s),$
for all $s\in\mathcal{S}$. They also share the same optimal action-value function, denoted $q_{*}$, which is defined as
@@
for all $s\in\mathcal{S}$. It holds that
@@
The optimal value functions satisfy the Bellman optimality equations, i.e.
$v_{*}(s)$ $=\max_{a\in\mathcal{A}}\mathbb{E}\bigl{[}R_{t+1}+\gamma v_{*}(S_{t+1})\bigm{|}S_{t}=s,A_{t}=a\bigr{]}$
$=\max_{a\in\mathcal{A}}\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\bigl{[}R_{a}(s,s^{\prime})+\gamma v_{*}(s^{\prime})\bigr{]}$
and
$q_{*}(s,a)$ $=\mathbb{E}\bigl{[}R_{t+1}+\gamma\max_{a^{\prime}\in\mathcal{A}}q_{*}(S_{t+1},a^{\prime})\bigm{|}S_{t}=s,A_{t}=a\bigr{]}$
$=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\bigl{[}R_{a}(s,s^{\prime})+\gamma\max_{a^{\prime}\in\mathcal{A}}q_{*}(s^{\prime},a^{\prime})\bigr{]},$
@@
For finite MDPs, the Bellman optimality equations have unique solutions independent of the policy. If the dynamics of the MDP in known, i.e. the $P$ and $R$ functions, one can solve the Bellman optimality equations. Once the optimal state-value function $v_{*}$ is known, any policy that is greedy with respect to $v_{*}$ is an optimal policy.
@@
Optimal policies can be computed with dynamic programming (DP) algorithms, such as policy iteration or value iteration, given a perfect model of the environment in the form of an MDP. DP algorithms are obtained by turning Bellman equations into update rules that improve the approximations of the desired value functions. However, the utility of DP algorithms is limited due to the perfect model assumption and its computational expense. Therefore, for large state space environments or cases there the model of the environment is unavailable, Monte Carlo methods or temporal-difference learning (TD) methods are used, see [31].
2.4 Q-learning
@@
Q-learning is a TD control algorithm and is one of the most popular model-free methods for solving RL problems. It is used to learn the optimal action-selection policy. The one-step Q-learning, which is the simplest form of Q-learning, is defined by the update rule
$Q(S_{t},A_{t})\leftarrow Q(S_{t},A_{t})+\alpha\bigl{[}R_{t+1}+\gamma\max_{a\in\mathcal{A}}Q(S_{t+1},a)-Q(S_{t},A_{t})\bigr{]}\!,$
@@
where $0<\alpha\leq 1$ is the learning rate, which determines how fast past experiences are forgotten by the agent, and the learned action-value function $Q$ directly approximates $q_{*}$. The algorithm can be outlined as follows. First, initialise $Q(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$ and start the environment at state $S_{0}$. For time steps $t=0,1,2,\ldots$, choose action $A_{t}$ according to the $\epsilon$-greedy policy, i.e.
@@
$A_{t}\leftarrow\begin{cases}\textrm{argmax}_{a\in\mathcal{A}}Q(S_{t},a)&\textrm{with probability }1-\epsilon\\ \textrm{uniformy random action in }\mathcal{A}&\textrm{with probability }\epsilon,\\ \end{cases}$
@@
and take action $A_{t}$ in the environment. In result, the state is changed from $S_{t}$ to $S_{t+1}\sim P_{a}(s,\cdot)$. Observe $S_{t+1}$ and $R_{t+1}$. Finally, revise the action-value function $Q$ at state-action $(S_{t},A_{t})$ in accordance with the above update rule.
@@
The Q-learning algorithm iteratively updates the Q-values based on observed transitions, gradually refining its estimates of the optimal action-value function. All that is required to assure convergence to $q_{*}$ is that all $(s,a)\in\mathcal{S}\times\mathcal{A}$ pairs continue to be updated. Notice that the convergence is guaranteed independently of the policy being followed. This property renders Q-learning an off-policy TD control algorithm and enables the use of the $\epsilon$-greedy policy, which is a simple technique for handling the exploration-exploitation tradeoff, to ensure the convergence. Once learned, the Q-values are used to provide the agent with the optimal policy for selecting actions in each state.
2.5 Deep reinforcement learning
@@
The Q-learning algorithm has proved successful for controlling small-scale environments. However, in the case of large state-action spaces, Q-learning cannot learn all action values in all states separately and can fail to converge to an optimal solution. To address this issue, the idea was conceived to learn a parameterised value function $Q(a,s;\theta)$ by updating the parameters $\theta$ in accordance with the standard Q-learning-derived rule:
$\theta_{t+1}\leftarrow\theta_{t}+\alpha\Big{[}R_{t+1}+\gamma\max_{a}Q(S_{t+1},a;\theta_{t})$
@@
resembling stochastic gradient descent. It was shown that as the agent explores the environment, the parameterized value function converges to the true $Q$-function, see, e.g., [31].
The above idea is realized in the deep RL (DRL) framework, which combines classical RL with artificial neural network (ANN) function approximation. In DRL, a function approximator is trained to estimate the Q-values. The approximator is a deep Q-network agent (DQN), i.e. a multi-layer ANN which for a given observed state of the environment $s$ outputs a vector of action values $Q(s,\cdot;\theta)$, where $\theta$ are the parameters of DQN. The aim of DQN is to approximate $q_{*}$ by iteratively minimising a sequence of loss functions $L_{t}(\theta_{t})$:
$L_{t}(\theta_{t})=\left(Y_{t}^{\textrm{DQN}}-Q(S_{t},A_{t};\theta_{t})\right)^{2},$
@@
$Y_{t}^{\textrm{DQN}}\triangleq R_{t+1}+\gamma\max_{a}Q(S_{t+1},a;\theta_{t}^{-}),$
@@
while $\theta_{t}^{-}$ and $\theta_{t}$ are the parameters of the target and online networks, respectively. The target network is the same as the on-line network, but its parameter values are copied from the online network every $k$ steps and kept fixed at other steps.
2.5.1 Double Deep Q-Network
@@
In the study of [18], the Double Deep Q-Network (DDQN) [34] was used to control BNs. Both standard Q-learning and DQN uses the same values both to select and to evaluate an action, which can lead to selection based on overestimated values [34]. DDQN combines Double Q-learning [34] with Deep Q-learning to address this issue and uses the following DDQN target In the study  [18], the Double Deep Q-Network (DDQN) from  [34] was used to control Boolean Networks. Both standard Q-learning and Deep Q-Network (DQN) use the same values for both selecting and evaluating actions, which can lead to selecting actions based on overestimated values, as noted by the authors. To address this issue, DDQN introduces Double Q-learning with separete networks for predicting actions and their values. It uses the following DDQN target:
@@
$Y_{t}^{\textrm{DDQN}}\triangleq R_{t+1}+\gamma Q(S_{t+1},\textrm{argmax}_{a}Q(S_{t+1},a;\theta_{t});\theta_{t}^{\prime}).$
The second set of parameters, $\theta_{t}^{\prime}$, enables fairly evaluation of the greedy policy. The DDQN parameters are updated by symmetrically switching the roles of $\theta_{t}$ and $\theta_{t}^{\prime}$.
2.5.2 Branching Dueling Q-Network
@@
In our study, to allow for simultaneous perturbation of a combination of genes within a DRL action, we replace the original DDQN architecture of [18] with the Branching Dueling Q-Network (BDQ) architecture [33]. BDQs are designed to address complex and high-dimensional action spaces; they enhance the scalability and sample efficiency of RL algorithms in complex scenarios. BDQs can be thought of as an adaptation of the dueling network architecture [36] into the action branching architecture. A BDQ uses two separate ANNs, i.e. the target network for evaluation and the controller network for selection of actions. Furthermore, instead of using a single output layer for all actions, BDQ introduces multiple branches, each responsible for a specific subset of actions. This decomposition allows for parallel processing and shared representations, making it more efficient for the agent to learn and execute tasks with a diverse action space. BDQs avoid overestimating Q-values, can more rapidly identify action redundancies, and generalise more efficiently by learning a general Q-value that is shared across many similar actions.
2.6 Graph Neural Networks
Graph Neural Networks (GNNs) have gained significant attention in recent years due to their ability to model and analyse graph-structured data. Graphs are powerful generic mathematical models that capture relationships and dependencies between entities in various domains, including social networks, biological systems, and recommendation engines. The main idea underlying GNNs is the use of pairwise message passing, i.e. graph nodes iteratively update their representations by exchanging information with their direct neighbours. In this study we focus on a specific type of GNNs, i.e. Graph Convolutional Networks.
2.6.1 Graph Convolutional Operation
@@
Graph Convolutional Networks (GCNs) are a subclass of GNNs and can be considered a generalization of classical convolutional neural networks (CNN) to graph-structured data. The essence of GCNs lies in their ability to perform graph convolutions, a key operation that allows them to learn and propagate information across the nodes of a graph. Graph convolution is inspired by traditional convolutional operations in image processing. While convolutional layers in traditional neural networks operate on grid-structured data, graph convolutional layers extend this concept to non-grid structures. The goal is to capture and aggregate information from neighbouring nodes, enabling the model to understand the relational structure within the graph. The information propagation happens through kernels – that is functions similar to kernels in CNNs but adapted to the graph’s structure.

### v2

@@ III-C Reinforcement learning
Nearly all RL algorithms involve estimating value functions. Herein, we focus on the action-value function $q_{\pi}(s,a)$ defined as the expected return starting from $s$, taking the action $a$, and thereafter following policy $\pi$:
$q_{\pi}(s,a)$ $\triangleq\mathbb{E}_{\pi}\left[\sum_{k=0}^{\infty}\gamma^{k}R_{t+k+1}\biggm{|}S_{t}=s,A_{t}=a\right]\!.$
The value function satisfies a particular recursive relationship expressed in the form of a Bellman equation, i.e.
$q_{\pi}(s,a)=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\Bigl{[}R_{a}(s,s^{\prime})+\gamma\sum_{a^{\prime}\in\mathcal{AC}}\pi(s^{\prime},a^{\prime})q_{\pi}(s^{\prime},a^{\prime})\Bigr{]}.$ (1)
Value functions define a partial order over policies and there always exists at least one policy which is better than or equal to all other policies [22]. This policy, denoted $\pi_{*}$, is referred to as the optimal policy. Optimal policies share the same optimal action-value function, denoted $q_{*}$, which is defined as
@@
for all $s\in\mathcal{S}$ and $a\in\mathcal{AC}$. It holds that
@@
The optimal action-value function satisfies the Bellman optimality equation:
$q_{*}(s,a)$ $=\mathbb{E}\bigl{[}R_{t+1}+\gamma\max_{a^{\prime}\in\mathcal{AC}}q_{*}(S_{t+1},a^{\prime})\bigm{|}S_{t}=s,A_{t}=a\bigr{]}$
$=\sum_{s^{\prime}\in\mathcal{S}}P_{a}(s,s^{\prime})\bigl{[}R_{a}(s,s^{\prime})+\gamma\max_{a^{\prime}\in\mathcal{AC}}q_{*}(s^{\prime},a^{\prime})\bigr{]},$
@@
Optimal policies can be computed with dynamic programming (DP) algorithms, such as policy iteration or value iteration, given a perfect model of the environment in the form of an MDP. DP algorithms are obtained by turning Bellman equations into update rules that improve the approximations of the desired value functions. However, the utility of DP algorithms is limited due to the assumption of a perfect model and its computational expense. Therefore, for large state space environments or cases where the model of the environment is unavailable, Monte Carlo methods or temporal-difference learning (TD) methods are used, see [22].
III-D Q-learning
@@
Q-learning is a TD control algorithm and is one of the most popular model-free methods to solve RL problems. It is used to learn the optimal action-selection policy. The one-step Q-learning is defined by the update rule
$Q(S_{t},A_{t})\leftarrow Q(S_{t},A_{t})+\alpha[\,R_{t+1}+\gamma\max_{a\in\mathcal{AC}}Q(S_{t+1},a)-Q(S_{t},A_{t})\!],$
@@
where $0<\alpha\leq 1$ is the learning rate, which determines how fast past experiences are forgotten by the agent, and the learned action-value function $Q$ directly approximates $q_{*}$. The algorithm can be outlined as follows. First, initialise $Q(s,a)$ for all $(s,a)\in\mathcal{S}\times\mathcal{AC}$ and start the environment at state $S_{0}$. For time steps $t=0,1,2,\ldots$, choose action $A_{t}$ according to the $\epsilon$-greedy policy, i.e.
@@
$A_{t}\leftarrow\begin{cases}\operatorname*{arg\,max}_{a\in\mathcal{AC}}Q(S_{t},a)&\textrm{with probability }1-\epsilon\\ \textrm{uniformy random action in }\mathcal{AC}&\textrm{with probability }\epsilon,\\ \end{cases}$
@@
and apply action $A_{t}$ in the environment. As a result, the state is changed from $S_{t}$ to $S_{t+1}\sim P_{a}(s,\cdot)$. Observe $S_{t+1}$ and $R_{t+1}$. Finally, revise the action-value function $Q$ at state-action $(S_{t},A_{t})$ in accordance with the above update rule.
@@
All that is required to assure convergence to $q_{*}$ is that all $(s,a)\in\mathcal{S}\times\mathcal{AC}$ pairs continue to be updated. Notice that the convergence is guaranteed independently of the policy being followed. This property renders Q-learning an off-policy TD control algorithm and enables the use of the $\epsilon$-greedy policy, which is a simple technique for handling the exploration-exploitation tradeoff, to ensure the convergence.
III-E Deep reinforcement learning
@@
The Q-learning algorithm has proven to be successful in controlling small-scale environments. However, in the case of large state-action spaces, Q-learning cannot learn all action values in all states separately and can fail to converge to an optimal solution. To address this issue, the idea was conceived to learn a parameterised value function $Q(s,a;\theta)$ by updating the parameters $\theta$ in accordance with the standard Q-learning-derived rule:
$\theta_{t+1}\leftarrow\theta_{t}+\alpha\Big{[}R_{t+1}+\gamma\max_{a\in\mathcal{AC}}Q(S_{t+1},a;\theta_{t})$
@@
resembling stochastic gradient descent. It was shown that as the agent explores the environment, the parameterized value function converges to the true $Q$-function, see, e.g. [22].
The above idea is realized in the deep RL (DRL) framework, which combines classical RL with artificial neural network (ANN) function approximation. In DRL, a function approximator is trained to estimate the Q-values. The approximator is a deep Q-network agent (DQN), i.e. a multi-layer ANN which for a given observed state of the environment $s$ outputs a vector of action values $Q(s,\cdot;\theta)$, where $\theta$ are the parameters of DQN. The aim of DQN is to approximate $q_{*}$ by iteratively minimising a sequence of loss functions $\mathcal{L}_{t}(\theta_{t})$:
$\mathcal{L}_{t}(\theta_{t})=\left(Y_{t}^{\textrm{DQN}}-Q(S_{t},A_{t};\theta_{t})\right)^{2},$
@@
$Y_{t}^{\textrm{DQN}}\triangleq R_{t+1}+\gamma\max_{a\in\mathcal{AC}}Q(S_{t+1},a;\theta_{t}^{-}),$
@@
while $\theta_{t}^{-}$ and $\theta_{t}$ are the parameters of the target and online networks, respectively. The target network is the same as the online network, but its parameter values are copied from the online network every $k$ steps and kept fixed at other steps.
III-E1 Double Deep Q-Network (DDQN)
@@
Both standard Q-learning and DQN use the same values for both selecting and evaluating actions, which can lead to choosing actions based on overestimated values [23]. To address this issue, DDQN introduces two networks with separate sets of weights, $\theta$ and $\theta^{\prime}$, where the former is used to predict actions and the latter to evaluate the greedy policy. They are trained by assigning each experience randomly to update one of the networks. DDQN uses the following DDQN target:
@@
$Y_{t}^{\textrm{DDQN}}\triangleq R_{t+1}+\gamma Q(S_{t+1},\operatorname*{arg\,max}_{a\in\mathcal{AC}}Q(S_{t+1},a;\theta_{t});\theta_{t}^{\prime}).$
III-E2 Branching Dueling Q-Network (BDQ)
@@
BDQs are designed to address complex and high-dimensional action spaces; they enhance the scalability and sample efficiency of RL algorithms in complex scenarios. BDQs can be thought of as an adaptation of the dueling network architecture [24] into the action branching architecture. A BDQ uses two separate ANNs, i.e. the target network $Q^{-}$ for evaluation and the controller network $Q$ for selection of actions. Furthermore, instead of using a single output layer for all actions, BDQ introduces multiple branches, each responsible for a specific subset of actions. This decomposition allows for parallel processing and shared representations, making it more efficient for the agent to learn and execute tasks with a diverse action space. BDQs avoid overestimating Q-values, can more rapidly identify action redundancies, and generalise more efficiently by learning a general Q-value that is shared across many similar actions.
The BDQ loss function is the mean squared temporal-difference error across the branches. Formally, the temporal-difference target is defined as:
$Y_{t}^{\textrm{BDQ(d)}}\triangleq R_{t+1}+\gamma Q_{d}^{-}(S_{t+1},\operatorname*{arg\,max}_{a\in\mathcal{AC}}Q_{d}(S_{t+1},a;\theta_{t});\theta_{t}^{-}),$
where $Q_{d}^{-}$ and $Q_{d}$ denote the branch $d$ of the target and controller network, respectively. Then the loss function is given by:
$\mathcal{L}_{t}^{\textrm{BDQ}}\triangleq\mathbb{E}_{(s,a,r,s^{\prime})\sim\mathcal{D}}\left[\frac{1}{N}\sum_{d}(Y_{t}^{\textrm{BDQ(d)}}-Q_{d}(s,a_{d};\theta_{t}))^{2}\right],$
where $a$ denotes the joint-action tuple $(a_{1},a_{2},\ldots,a_{n})$ and $\mathcal{D}$ is a (prioritised) experience replay buffer.
III-F Graph Neural Networks
@@
Graph Neural Networks (GNNs) have gained significant attention in recent years due to their ability to model and analyse graph-structured data. Graphs are powerful generic mathematical objects that capture relationships and dependencies between entities in various domains, including social networks, biological systems, and recommendation engines. The main idea underlying GNNs is the use of pairwise message passing, i.e. graph nodes iteratively update their representations by exchanging information with their direct neighbours. In this study, we focus on a specific type of GNNs, i.e. Graph Convolutional Networks (GCNs). GCNs are a subclass of GNNs and can be considered a generalization of classical convolutional neural networks (CNN) to graph-structured data. The essence of GCNs lies in their ability to perform graph convolution, a key operation that allows them to learn and propagate information across the nodes of a graph. Graph convolution is inspired by traditional convolutional operations in image processing. While convolutional layers in traditional neural networks operate on grid-structured data, graph convolutional layers extend this concept to non-grid structures. The goal is to capture and aggregate information from neighbouring nodes, enabling the model to understand the relational structure within the graph. The information propagation happens through kernels – that is functions similar to kernels in CNNs, but adapted to the graph’s structure.
IV Control problem formulation
We proceed to formulate a novel target-control problem tailored for cellular reprogramming, understood as an artificial change of the identity or functional state of a cell by manipulating its gene expression profile.
Often biologists are interested not in reaching a single, fully specified state, but rather a set of states in which a specified subset of genes attains a certain configuration of expression levels, which remains constant afterward. For example, they may be interested in achieving any state in which apoptosis occurs. Therefore, we assume that the target is given as a combination of desired values for a subset of genes rather than a fully specified attractor as in [7]. We define an attractor-target control strategy as follows.
[…]
Figure 1: Schematic illustration of the concept behind Attractor-Target Control. The source attractor $\mathcal{A}_{S}$ (blue dot) is not aligned with the target configuration (target gene expression profile for two of the four genes is encoded with colours: olivegreen - OFF, aquamarine - ON). The target can be any of the attractors (red dots) containing a state that is aligned with the target configuration. Transitions that follow the original network dynamics are shown with dashed arrows. Red lightnings indicate interventions (control actions) and red arrows indicate the change of state due to interventions. The depicted control strategy is of length 3, which from $\mathcal{A}_{S}$ traverses two intermediate attractor states (green dots) to finally reach one of the target attractors. Ellipses represent basins of attraction of the respective attractors on the control trajectory (path).
Definition IV.1 (Attractor-Target Control Strategy).
Given a BN, a source (pseudo-)attractor state, and a specified target configuration for a subset of genes, an attractor-target control strategy is a sequence of instantaneous interventions (perturbations) which guides the BN dynamics from the source (pseudo-)attractor state to some (pseudo-)attractor state that aligns with the desired gene configuration. An instantaneous intervention is understood as a simultaneous flip of values for a subset of genes in a particular network state that is being perturbed. This results in repositioning the dynamics to a new state in the state space of the BN/PBN. Instantaneous means that directly after the relocation in the state space, the BN/PBN evolves in accordance with its original dynamics starting from the new state. The application of interventions is limited to (pseudo-)attractor states.
Furthermore, the length of a control strategy is defined as the number of interventions in the control sequence. We refer to a target control strategy of the shortest length as the minimal attractor-target control strategy.
The concept of the attractor-target control strategy is schematically illustrated in Figure 1 and an example is provided in Appendix -A. Notice that, according to the above definition, we aim to guide the network dynamics toward any state within a subset of (pseudo-)attractor states that match the target configuration, rather than a specific (pseudo-)attractor. This definition is more general than the definition of the ‘Attractor-Based Control Strategy’ introduced in [7], where the system is expected to reach a specific attractor. The current definition of the ‘Attractor-Target Control Strategy’ aligns better with contemporary biological practice.
Definition IV.2 (Attractor-Target Control).
V GATTACA: Graph-based Attractor-Target Control Algorithm
We propose a Graph-based Attractor-Target Control Algorithm (GATTACA) computational framework – a GNN-based DRL approach for solving the attractor-target control problem. With cellular reprogramming in mind, we focus on cellular phenotypic functional states, which are observable in experimental practice, and thus allow the DRL agent, which we refer to as the Attractor-Target Graph-based Controller, to take actions only in (pseudo-)attractor states (PA states for short) in agreement with our control problem formulation in Section IV. Since the identification of all attractor states is an NP-hard problem in itself, we improve the pseudo-attractor states identification procedure (PASIP), which was previously established in [7]. The improved PASIP and the Attractor-Target Graph-based Controller constitute the two main components of the GATTACA framework. The source code of our framework together with the trained models are available at [26].
V-A Improved PASIP (iPASIP)
We observed that the original PASIP method in [7] can be impeded by the presence of large attractors in a BN model. For example, the T-LGL network of [27] has an attractor with more than 1000 states. Of course, if the existence of a large attractor is foreseen and a rough estimate of its size is available, the PASIP $k_{1}$ and $k_{2}$ hyperparameters could be adjusted appropriately based on Theorem .1 presented in Appendix -C, which was originally proved in [7]. However, this would imply the necessity of longer simulations of the BN model, which are computationally expensive. If the existence of a large attractor is not anticipated and the hyper-parameters are not adjusted, the original PASIP is prone to getting trapped in such an attractor while waiting for one of the attractor states to demonstrate a pronounced frequency of re-visits. To handle such cases effectively, we modify PASIP by adding additional checkpoints that enable the identification of a representative state within a large attractor. The improved PASIP (iPASIP) consists of steps executed in two phases: (I) environment pre-processing and (II) DRL agent training. It extends the original PASIP from [7] by incorporating Steps I-2 and II-3 as additional checkpoints.
  1. I-1
    During environment pre-processing, $k$ trajectories are simulated, each starting from a randomly selected initial state. Each BN simulation is run for initial $n_{0}=200$ time steps, which are discarded, i.e. the so-called burn-in period. Then, the simulation continues for $n_{1}=1,000$ time steps during which the visits to individual states are counted. All states in which at least $k_{1}=5\%$ of the simulation time $n_{1}$ is spent are added to the list of PA states.
  2. I-2
    If at any point more than $d_{1}=1,000,000$ steps are taken without encountering a dominating state (one revisited at least $k_{1}=5\%$ of the time), it is assumed that the dynamics have become trapped in a large attractor. The most frequently encountered state is then added to the list of PA states as a representative of this plausible attractor.
  3. II-1
    The simulation of the BN environment may enter a fix-point attractor not detected in Step I. If the simulation gets stuck in a particular state for $n_{2}=1000$ steps, the state is verified to be a fixed-point attractor and added to the list of PA states.
  4. II-2
    During training, the simulation of the BN environment may enter a multi-state attractor that was not detected in Step I. For this reason, a history of the most recently visited states is maintained. When the history buffer reaches the size of $n_{3}=10,000$ items, revisits for each state are counted and states revisited more than $k_{2}=15\%$ of times are added to the list of PA states. If no such state exists, the history buffer is cleared and the procedure continues for another $n_{3}$ time steps. The new PA states are added provided no known PA state was reached. Otherwise, the history information is discarded.
  5. II-3
    As in I-2, if at any point more than $d_{2}=1,000,000$ steps are taken without encountering a dominating state (one revisited at least $k_{1}=5\%$ of the time), it is assumed that the dynamics have become trapped in a large attractor and a representative of this plausible attractor is added to the list of PA states.
iPASIP features a number of hyperparameters, i.e. $n_{0}$, $n_{1}$, $n_{2}$, $n_{3}$, $d_{1}$, $d_{2}$, $k_{1}$, and $k_{2}$. Their values should be set on a case-by-case basis in accordance with prior knowledge about the system under study, or some preliminary experiments, and Theorem .1 in Appendix -C. The values provided above were suitable for our evaluations discussed in Section VI.
We notice that while the new checkpoints may clearly introduce false-positive PA states, this occurs rarely. Importantly, the representative states of plausible large attractors stand out due to being frequently revisited. Even if these states are not exact PA states, to a large extent they resemble them. Therefore, given their rarity, adding them to the list of PA states is acceptable also from the biological point of view. This is confirmed by the results of our case studies, presented and discussed in Section VII.
Since iPASIP can introduce spurious PA states, we shall henceforth use the term ‘PA state’ in a broader sense: it will refer to all states returned by iPASIP rather than only those defined in Definition III.1.
V-B Attractor-Target Graph-based Controller (ATGC)
Our DRL agent, which we refer to as the Attractor-Target Graph-based Controller (ATGC), consists of three main components: a GCN, a multi-layer perceptron (MLP), and a BDQ network. The architecture of ATGC is schematically illustrated in Figure 2.